OpenStudy (anonymous):

PLEASE HELP REALLY QUICK!! How many liters of chlorine gas can react with 56.0 grams of calcium metal at standard temperature and pressure? Ca + Cl2 ----> CaCl2 How go I go about solving this? Where do I even start?

OpenStudy (aaronq):

1. balance equation (already balanced) 2. convert the mass (m) given to moles (n). Hint: use molar mass (M). \[n=\frac{ m }{ M }\] Now, the moles of Cl2 (g) and Ca(s) should be the same, so they can react completely (i.e. nothing is left).. theoretically, so... 3. write an expression relating the 2 species. use the coefficients and the moles (n) of each species in a ratio, like this: For a general equation: aA +bB -> cC where a,b,c (lowercase) are the coefficients and A,B,C are the species \[\frac{ n _{A} }{ a }=\frac{ n _{B} }{ b }=\frac{ n _{C}}{ c }\]: Do the same for the equation you have.